please kindly assist me with corrections for my calculations.
Determine the amount that has to be invested now, to be worth R10 000 in 8 months’ time, if the annual
simple interest rate is 9,75%.
formula P = I/RT
P = ?
I = 10 000
R = 9.75
T = 8/12
: 10 000/9.75 x 8/12
: 1538.46 which is suppose to be the initial
to confirm my answer l will say : 1538.46 x 9.75 x 8/12 = 999.99
options given
[1] R9 389,67
[2] R9 398,61
[3] R1 333,33
[4] R6 060,61

Answer by Alan3354(69443)

Determine the amount that has to be invested now, to be worth R10 000 in 8 months’ time, if the annual
simple interest rate is 9,75%.
formula P = I/RT
-------------------
8 months = 2/3 years
9.75%/yr --> 0.0975
AMT = starting amount
-----
Interest = AMT*(2/3)*0.0975 = AMT*0.065
---
AMT + AMT*0.065 = 10000
AMT = 10000/(1.065)
AMT = 9389.67